# HV-BEV: Decoupling Horizontal and Vertical Feature Sampling for Multi-View 3D Object Detection

### v2

@@ III-D Horizontal Cross-View Feature Aggregation
The standard spatial cross-attention (SCA) mechanism projects each 3D reference point to all image coordinate systems, summing features extracted from all hit views; this approach partially aggregates cross-view spatial information for 3D points that can be projected into adjacent views after coordinate transformation. However, for reference points projected into only one view, the imaging information of their associated objects may span multiple views, leading to inadequate feature aggregation. In this section, we present a graph-based horizontal cross-view feature aggregation (HCA) module, which learns a graph structure on the horizontal plane for each 3D reference point. As illustrated in Fig.4, this module refines each point’s feature representation by integrating sampled features from neighboring nodes, thereby enhancing spatial correlations and improving global feature aggregation.
@@
Formally, for any 3D reference point $p_{3d}\in(x,y,z)$ in $Q_{p}$, where $(x,y)$ represents the horizontal center coordinates of the corresponding grid region and $z$ is the vertical height coordinate, we define a graph structure $G=(\mathbf{A},\mathbf{X},\mathbf{V})$ centered on $p_{3d}$. Here, $\mathbf{V}=\{v_{k}\}_{k=1}^{M}$ represents the set of nodes, where $M$ denotes the number of $p_{3d}$ and its neighbors, and $v_{1}$ is the point $p_{3d}$ itself; The adjacency matrix $\mathbf{A}\in\mathbb{R}^{M\times M}$ is defined as a symmetric binary matrix representing an undirected graph, ensuring bidirectional connectivity between $p_{3d}$ and its neighbors; And $\mathbf{X}\in\mathbb{R}^{M\times C}$ represents the feature embeddings of all nodes, where the $m$-th feature vector $\mathbf{x}_{k}\in\mathbb{R}^{C}$ corresponds to the attribute of $v_{k}$. Furthermore, each neighboring node $v_{k}(k=2,\dots,M)$ is learned from a linear layer:
$\begin{array}[]{l}v_{k}=p_{3d}+\bigtriangleup p_{xy}=(x+\bigtriangleup x,y+\bigtriangleup y,z)\end{array},$ (5)
$\begin{array}[]{l}\bigtriangleup p_{xy}=\text{Linear}(Q_{p})\in\mathbb{R}^{2}\end{array},$ (6)
@@
where $\bigtriangleup p_{xy}=(\bigtriangleup x,\bigtriangleup y)$ represent the sampling offsets along two dimensions of the horizontal plane. The feature vector of $p_{3d}$, denoted as $\mathbf{x}_{1}$, is obtained via deformable attention, similar to the vanilla SCA:
@@
$\begin{array}[]{l}\mathbf{x}_{1}=\dfrac{1}{|\mathcal{V}_{hit}|}\sum\limits_{i\in\mathcal{V}_{\text{hit}}}\text{DefAttn}(Q_{p},\mathcal{P}(p_{3d},i),F_{t}^{i})\end{array},$ (7)
@@
To optimize computational efficiency, the features of other neighboring nodes $v_{k}(k=2,\dots,M)$ are extracted from the corresponding image features using bilinear interpolation:
@@
After obtaining the attributes of all nodes, we update the centered node features via:
$\begin{array}[]{l}\mathbf{x}^{\prime}_{1}=\mathbf{x}_{1}+\sum\limits_{k=2}^{M}\mathbf{A}_{1k}\cdot\mathbf{x}_{k}\cdot\mathbf{w}_{k-1}\end{array}.$ (11)
Here, $\mathbf{w}$ represents the edge weight vector between $p_{3d}$ and its neighbors, learned through a linear layer followed by a softmax function. Finally, the features of all $N_{ref}$ reference points are aggregated via a weighted sum, yielding the refined query $Q_{p}$ as the output of HCA module, fully enriched with spatial information.
Figure 4: The process of horizontal cross-view feature aggregation module. Each 3D reference point constructs a graph structure on its corresponding horizontal plane, where all neighboring nodes are projected onto the respective camera views using intrinsic parameters to extract features. These features are then aggregated through learned edge weights, effectively weighting the contributions of each neighboring node. This process enables the comprehensive aggregation of 3D spatial features for each query.
@@ III-E Overall Loss
where $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ are the balance weights assigned to the respective loss terms. We adopt focal loss [46] and L1 loss as the loss functions for classification and regression, respectively. $L_{hgt}$ is a cross-entropy loss function defined to provide deep supervision for the estimated height distribution in the vertical adaptive height-aware sampling module. During training, we apply sparse supervision to $H^{fus}$ through utilizing the height of the annotated 3D bounding box center.
Specifically, to facilitate the calculation of the discrepancy between the ground-truth height and the discretized predicted height distribution, we first derive the discretized ground-truth height distribution $H_{p}^{gt}$ corresponding to each BEV grid region. Since an object’s height is inherently a range rather than a fixed value, we employ a Gaussian kernel $K$ to compute the differences between the annotated center height of an object and the predefined height bins. Height bins closer to the center are assigned higher probabilities. For convenience, we use an indicator $I_{p}$ to denote whether the cell grid centered at $p$ falls within the object’s region. The probability corresponding to the $m$-th height bin of $H_{p}^{gt}$ is calculated as:
where $z_{gt}$ denotes the height of ground truth center, $z_{m}$ represents the center value of the $m$-th height bin, and $\sigma$ is the standard deviation controlling the spread of the Gaussian kernel. We then calculate the height distribution loss using cross-entropy as follows:
Figure 5: Visualization results of HV-BEV and the baseline. The green arrows highlight instances where our method achieves improved detection performance on large cross-view objects. In BEVFormer, such objects are often missed or misdetected due to view truncation. However, our approach enhances detection by leveraging a global focus on the structured characteristics of objects, effectively mitigating these challenges.
@@ IV-A Datasets and Metrics
Dataset We train and validate our model on the publicly available nuScenes [47] autonomous driving dataset. This dataset is constructed using six cameras providing a full $360^{\circ}$ field of view, with images captured at a resolution of 1600$\times$900 and a frequency of 12Hz. It also incorporates LiDAR sensors operating at 20Hz with a $360^{\circ}$ horizontal field of view. The complete dataset consists of 1,000 diverse and challenging autonomous driving video sequences, collected from four distinct locations under varying weather and lighting conditions. Each sequence spans 20 seconds. Keyframes were sampled at a frequency of 2Hz for annotation, with 3D bounding boxes labeled for 10 object categories. Following the official dataset split, 28,130 frames from 750 sequences are used for training, while 6,019 and 6,008 frames from the remaining 300 sequences are designated for validation and testing, respectively.
Metrics We evaluate the performance of our model using the 3D object detection metrics provided by the official nuScenes benchmark. These metrics include mean Average Precision (mAP), mean Average Translation Error (mATE), mean Average Scale Error (mASE), mean Average Orientation Error (mAOE), mean Average Velocity Error (mAVE), mean Average Attribute Error (mAAE), and the nuScenes Detection Score (NDS). Among these, mAP and NDS are the primary evaluation metrics.
@@ IV-B Implementation Details
Our model adopts BEVFormer as the baseline, with configurations tailored for two settings: $base$ and $small$. Both settings share the same backbone, ResNet101-DCN [44, 48] and VoVnet-99[54], initialized using pretrained parameters from FCOS3D [49] and DD3D[55], respectively. In the $base$ setting, input images have a resolution of $1600\times 900$, producing multi-scale features across four levels from the FPN, with a BEV grid size of $200\times 200$, a historical sequence length of 4, and a BEV encoder comprising 6 layers. In contrast, the $small$ setting uses input images of $1280\times 720$, generates single-scale feature maps, and adopts a BEV grid size of $150\times 150$, a historical sequence length of 3, and a BEV encoder with 3 layers. This design balances computational cost and performance while maintaining alignment with BEVFormer’s core architecture.
For the vertical adaptive height-aware module, the number of height bins is $D=8$, and the number of 3D reference points sampled in each query is $N_{ref}=4$. In the HCA module, each 3D reference point considers four neighboring nodes, and two sampling points are used for deformable attention when extracting features projected onto each 2D view. During training, the model uses the AdamW [50] optimizer with a cosine annealing learning rate schedule [51]. The initial learning rate is set to $2\times 10^{-4}$ and the weight decay is 0.01. Training is conducted on 8 NVIDIA L20 GPUs for 24 epochs.
@@ IV-C Main Results
Performance Improvements Over Baseline To comprehensively evaluate the performance improvements of our model, we configured it based on the three official configuration files provided by BEVFormer (i.e., tiny, small, and base) and trained it for the same number of epochs. The performance comparison on the nuScenes validation set is shown in Tab. I. Our model consistently outperforms the baseline in both key metrics, mAP and NDS, highlighting the effectiveness of our proposed optimization strategies. Under the $base$ configuration, our model achieved improvements of 2.3% in mAP and 1.6% in NDS compared to the baseline. In Fig. 5, we visualize a comparison of 3D detection results between the proposed method and the baseline on a sample of nuScenes validation set. For large objects near the ego vehicle, our method demonstrates a significant improvement in detection accuracy. Fig. 6 illustrates the bird’s-eye view detection results for the same sample in the LIDAR point cloud. It can be observed that our model exhibits fewer false positives and achieves a higher degree of overlap with the ground truth boxes. However, small distant objects remain challenging and are more prone to missed detections. Fig. 7 presents several robust detection results of HV-BEV for cross-view objects or large-scale objects. Our model demonstrates exceptional performance in handling such challenging scenarios.
Figure 6: Comparison of detection results in the LiDAR point cloud under the BEV perspective. (a) Baseline; (b) Ours. Green boxes represent ground truth annotations, while blue boxes denote predicted bboxes.
@@
BEVFormer tiny $50\times 50$ ResNet50 0.252 0.354
HV-BEV(Ours) tiny $50\times 50$ ResNet50 0.284 0.378
BEVFormer small $150\times 150$ ResNet101 0.370 0.479
HV-BEV(Ours) small $150\times 150$ ResNet101 0.399 0.502
BEVFormer base $200\times 200$ ResNet101 0.416 0.517
HV-BEV(Ours) base $200\times 200$ ResNet101 0.439 0.533
Comparison with State-of-the-Arts In addition to comparing our model with the baseline, we also present a performance comparison between our approach and some other state-of-the-art 3D odject detection methods on the nuScenes validation and testing sets. The results are summarized in Tab. II and Tab. III, respectively. On the validation set, we compared the performance of two model configurations. The $small$ configuration achieved 39.9% mAP and 50.2 NDS, surpassing one of the latest methods, QAF2D[52], under similar settings. The $base$ configuration achieved 43.9% mAP and 53.3% NDS, outperforming HeightFormer, which also incorporates explicit height information, under comparable configurations. On the testing set, we evaluated our method using two mainstream backbone networks. With VoVNet-99 as the backbone, our approach achieved 50.5% mAP and 59.8% NDS. Notably, our model demonstrated exceptional performance on the mAAE metric, likely due to its comprehensive consideration of spatial correlations of an object.
TABLE II: 3D detection results on nuScenes $val$ set. “C” indicates Camera. *: Trained with Lidar data. †: Pre-trained from FCOS3D. ‡: Pre-trained from DD3D.
@@
TABLE III: 3D detection results on nuScenes $test$ set. “C” and ”L” indicates Camera and LIDAR. *: Trained with Lidar data. †: Pre-trained from FCOS3D. ‡: Pre-trained from DD3D.
@@
CenterPoint[56] L - - 0.580 0.655 - - - - -
PointPainting[57] L+C - - 0.464 0.581 0.388 0.271 0.496 0.247 0.111
BEVFusion[58] L+C - - 0.717 0.738 0.252 0.238 0.321 0.265 0.126
BEVDet4D[14] C $1600\times 640$ Swin-T 0.451 0.569 0.511 0.241 0.386 0.301 0.121
GraphDETR3D†[23] C $1600\times 640$ ResNet101 0.418 0.472 0.668 0.250 0.440 0.876 0.139
PETRv2†[22] C $1600\times 640$ ResNet101 0.456 0.553 0.601 0.249 0.391 0.382 0.123
HV-BEVbase†(Ours) C $1600\times 900$ ResNet101 0.467 0.562 0.586 0.261 0.362 0.362 0.113
DETR3D‡[17] C $1600\times 900$ VoVNet-99 0.412 0.479 0.641 0.255 0.394 0.845 0.133
BEVDepth*‡[15] C $1408\times 512$ VoVNet-99 0.503 0.600 0.445 0.245 0.378 0.320 0.126
GraphDETR3D‡[23] C $1600\times 640$ VoVNet-99 0.425 0.495 0.621 0.251 0.386 0.790 0.128
PETRv2‡[22] C $1600\times 640$ VoVNet-99 0.490 0.582 0.561 0.243 0.361 0.343 0.120
VEDet‡[59] C $1600\times 640$ VoVNet-99 0.505 0.585 0.545 0.244 0.346 0.421 0.123
HeightFormer‡[40] C $1600\times 900$ VoVNet-99 0.481 0.573 - - - - -
HV-BEVbase‡(Ours) C $1600\times 900$ VoVNet-99 0.505 0.598 0.544 0.249 0.353 0.318 0.107
Figure 7: Some robust detection results for cross-view objects and large-scale objects (such as trucks) on the nuScenes validation set. Different colors represent different object classes.
IV-D Ablation Study

### v3

@@ III-D Dynamic Horizontal Cross-View Feature Aggregation
The standard spatial cross-attention (SCA) mechanism projects each 3D reference point to all image coordinate systems, summing features extracted from all hit views; this approach partially aggregates cross-view spatial information for 3D points that can be projected into adjacent views after coordinate transformation. However, for reference points projected into only one view, the imaging information of their associated objects may span multiple views, leading to inadequate feature aggregation. In this section, we present a dynamic horizontal cross-view neighborhood aggregation (DHCA) module, which learns a set of relevant neighboring points on the horizontal plane for each 3D reference point, forming a graph-like structure centered around the reference point. For clarity of presentation, we utilize the adjacency matrix and edges in a graph structure to represent the relationships between each reference point and its neighboring points. As illustrated in Fig.4, this module refines each point’s feature representation by integrating sampled features from neighboring points, thereby enhancing spatial correlations and improving global feature aggregation.
@@
Formally, for any 3D reference point $p_{3d}\in(x,y,z)$ in $Q_{p}$, where $(x,y)$ represents the horizontal center coordinates of the corresponding grid region and $z$ is the vertical height coordinate, we define a set of neighboring points $\mathbf{V}=\{v_{k}\}_{k=1}^{M}$ centered on $p_{3d}$, where $M$ denotes the number of $p_{3d}$’s neighbors. And $\mathbf{X}\in\mathbb{R}^{M\times C}$ represents the feature embeddings of all neighboring points, where the $k$-th feature vector $\mathbf{x}_{k}\in\mathbb{R}^{C}$ corresponds to the attribute of $v_{k}$. Furthermore, each neighboring point $v_{k}(k=1,\dots,M)$ is learned from a linear layer:
$\begin{array}[]{l}v_{k}=p_{3d}+\bigtriangleup p_{xy}^{k}=(x+\bigtriangleup x_{k},y+\bigtriangleup y_{k},z)\end{array},$ (9)
$\begin{array}[]{l}\bigtriangleup p_{xy}^{k}=\text{Linear}(Q_{p})\in\mathbb{R}^{2}\end{array},$ (10)
@@
where $\bigtriangleup p_{xy}^{k}=(\bigtriangleup x_{k},\bigtriangleup y_{k})$ represent the sampling offsets along two dimensions of the horizontal plane. Notably, since the distribution of associated features may vary across different heights of an object, the sampling offsets for different reference points in $Q_{p}$ are obtained individually through the shared linear layer. The feature vector of $p_{3d}$, denoted as $\mathbf{x}_{p}$, is obtained via deformable attention, similar to the vanilla SCA:
@@
$\begin{array}[]{l}\mathbf{x}_{p}=\dfrac{1}{|\mathcal{V}_{hit}|}\sum\limits_{i\in\mathcal{V}_{\text{hit}}}\text{DefAttn}(Q_{p},\mathcal{P}(p_{3d},i),F_{t}^{i})\end{array},$ (11)
@@
where $(x_{i},y_{i},z_{i})$ denotes the pixel $(x_{i},y_{i})$ with depth $z_{i}$ projected onto the $i$-th camera image.
To optimize computational efficiency, the features of the neighboring points $v_{k}(k=1,\dots,M)$ are extracted from the corresponding image features using bilinear interpolation:
@@
$\mathcal{P}(v_{k},i)$ represents the projection of $v_{k}$ onto the $i$-th camera view using the camera intrinsic matrix, and $\mathcal{V}_{hit}$ denotes the hit views.
After obtaining the attributes of all neighboring points, we update the centered reference point feature via:
$\begin{array}[]{l}\mathbf{x}^{\prime}_{p}=\mathbf{x}_{p}+\sum\limits_{k=1}^{M}\mathbf{w}_{k}\cdot\mathbf{x}_{k}\end{array}.$ (15)
Here, $\mathbf{w}_{k}$ represents the edge weight vector between $p_{3d}$ and its neighbors, learned through a linear layer followed by a softmax function. Finally, the features of all $N_{ref}$ reference points are aggregated via a sum operation, yielding the refined query $Q_{p}$ as the output of DHCA module, fully enriched with spatial information.
Figure 4: The process of the dynamic horizontal cross-view feature aggregation module. Each 3D reference point constructs a set of neighboring points on its corresponding horizontal plane, where all points are projected onto the respective camera views using intrinsic parameters to extract features. These features are then aggregated through learned edge weights, effectively weighting the contributions of each neighbors. This process enables the comprehensive aggregation of 3D spatial features for each query.
@@ III-E Overall Loss
where $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ are the balance weights assigned to the respective loss terms. We adopt focal loss [43] and L1 loss as the loss functions for classification and regression, respectively. $L_{hgt}$ is a cross-entropy loss function defined to provide deep supervision for the estimated height distribution in the vertical adaptive height-aware sampling module (See Eq. 8).
(a) Case 1
(b) Case 2
Figure 5: Visualization results of HV-BEV and the baseline on nuScenes validation set. Different colors of 3D bounding boxes denote different object classes. The green arrows highlight instances where our method achieves improved detection performance on large objects. In BEVFormer, such objects are often missed or misdetected due to view truncation or their occupancy across multiple grid cells. However, our approach enhances detection by leveraging a global focus on the structured characteristics of objects, effectively mitigating these challenges. On the right are the BEV-view detection results in LiDAR point clouds, with the top showing the results of BEVFormer and the bottom showing the results of HV-BEV. Green boxes represent ground truth annotations, while blue boxes denote predicted bboxes.
@@ IV-A Datasets and Metrics
The nuScenes Dataset We train and validate our model on the publicly available nuScenes [44] autonomous driving dataset. This dataset is constructed using six cameras providing a full $360^{\circ}$ field of view, with images captured at a resolution of 1600$\times$900 and a frequency of 12Hz. It also incorporates LiDAR sensors operating at 20Hz with a $360^{\circ}$ horizontal field of view. The complete dataset consists of 1,000 diverse and challenging autonomous driving video sequences, collected from four distinct locations under varying weather and lighting conditions. Each sequence spans 20 seconds. Keyframes were sampled at a frequency of 2Hz for annotation, with 3D bounding boxes labeled for 10 object categories. Following the official dataset split, 28,130 frames from 750 sequences are used for training, while 6,019 and 6,008 frames from the remaining 300 sequences are designated for validation and testing, respectively. We evaluate the performance of our model using the 3D object detection metrics provided by the official nuScenes benchmark. These metrics include mean Average Precision (mAP), mean Average Translation Error (mATE), mean Average Scale Error (mASE), mean Average Orientation Error (mAOE), mean Average Velocity Error (mAVE), mean Average Attribute Error (mAAE), and the nuScenes Detection Score (NDS). Among these, mAP and NDS are the primary evaluation metrics.
The Lyft Dataset To evaluate the generalizability of the proposed method, we further trained and validated its performance on the Lyft Level 5 AV dataset [54]. This dataset comprises data collected using seven cameras and three LiDAR sensors, with over 55,000 manually annotated 3D bounding boxes. The data format is similar to that of the nuScenes dataset. We utilized images from six camera views—front, front-left, front-right, back, back-left, and back-right—which are consistent with those used in nuScenes, as input to our model. A total of 22,680 frames were used for training, and 27,468 frames for testing. Since the Lyft dataset does not provide an official split between training and validation sets, we followed the split scheme recommended by MMDetection3D [55], using 18,900 frames for training and the remaining 3,780 frames for validation. For evaluation, we adopted a COCO-style mean Average Precision (mAP) metric, computing mAP over a range of 3D Intersection over Union (IoU) thresholds from 0.5 to 0.95. It is worth noting that an IoU threshold greater than 0.7 constitutes a stringent criterion for 3D object detection methods, which can result in relatively lower overall performance scores.
@@ IV-B Implementation Details
Our model adopts BEVFormer as the baseline, with configurations tailored for three settings: $base$, $small$ and $tiny$. All settings share the same backbone, ResNet101-DCN [41, 45] and VoVnet-99[51], initialized using pretrained parameters from FCOS3D [46] and DD3D[52], respectively. In the $base$ setting, input images have a resolution of $1600\times 900$, producing multi-scale features across four levels from the FPN, with a BEV grid size of $200\times 200$, a historical sequence length of 4, and a BEV encoder comprising 6 layers. In contrast, the $small$ setting uses input images of $1280\times 720$, generates single-scale feature maps, and adopts a BEV grid size of $150\times 150$, a historical sequence length of 3, and a BEV encoder with 3 layers. This design balances computational cost and performance while maintaining alignment with BEVFormer’s core architecture. The $tiny$ configuration retains the same settings as $small$, except for a further reduced BEV grid size of 50×50.
For the vertical adaptive height-aware module, the number of height bins is $D=8$, and the number of 3D reference points sampled in each query is $N_{ref}=4$. The standard deviation of the Gaussian kernel used to compute the difference between the annotated box center height and the height bin values is set to $\sigma=1$. In the DHCA module, each 3D reference point considers four neighboring nodes, and two sampling points are used for deformable attention when extracting features projected onto each 2D view. During training, the model uses the AdamW [47] optimizer with a cosine annealing learning rate schedule [48]. The initial learning rate is set to $2\times 10^{-4}$ and the weight decay is 0.01. Training is conducted on 8 NVIDIA L20 GPUs for 24 epochs.
@@ IV-C Main Results
Performance Improvements Over Baseline To comprehensively evaluate the performance improvements of our model, we configured it based on the three official configuration files provided by BEVFormer (i.e., tiny, small, and base) and trained it for the same number of epochs. The performance comparison on the nuScenes validation set is shown in Tab. I. Our model consistently outperforms the baseline in both key metrics, mAP and NDS, highlighting the effectiveness of our proposed optimization strategies. Under the $base$ configuration, our model achieved improvements of 2.3% in mAP and 1.6% in NDS compared to the baseline. The FPS evaluation was consistently performed on a single NVIDIA RTX 4090 GPU. While the proposed method introduces a marginal increase in inference time over the baseline, the resulting gain in detection accuracy comes at a negligible cost to efficiency. In Fig. 5, we visualize a comparison of 3D detection results between the proposed method and the baseline on a sample of nuScenes validation set. For large objects near the ego vehicle, our method demonstrates a significant improvement in detection accuracy. The right illustrates the bird’s-eye view detection results for the same sample in the LIDAR point cloud. It can be observed that our model exhibits fewer false positives and achieves a higher degree of overlap with the ground truth boxes. However, small distant objects remain challenging and are more prone to missed detections. This is primarily due to the limited visual information available for small and distant objects in the image, which leads to weak feature representations. The downsampling process in the feature extraction pipeline further degrades these representations, making it one of the key bottlenecks in vision-based 3D object detection.
@@
BEVFormer tiny $50\times 50$ ResNet50 0.252 0.354 29.75
HV-BEV(Ours) tiny $50\times 50$ ResNet50 0.284 0.378 27.91
BEVFormer small $150\times 150$ ResNet101 0.370 0.479 11.53
HV-BEV(Ours) small $150\times 150$ ResNet101 0.399 0.502 9.08
BEVFormer base $200\times 200$ ResNet101 0.416 0.517 5.83
HV-BEV(Ours) base $200\times 200$ ResNet101 0.439 0.533 4.38
Comparison with State-of-the-Arts In addition to comparing our model with the baseline, we also present a performance comparison between our approach and some other state-of-the-art 3D odject detection methods on the nuScenes validation and testing sets. The results are summarized in Tab. II and Tab. III, respectively. On the validation set, we compared the performance of two model configurations. The $small$ configuration achieved 39.9% mAP and 50.2% NDS, surpassing one of the latest methods, QAF2D[49], under similar settings. The $base$ configuration achieved 43.9% mAP and 53.3% NDS, outperforming HeightFormer, which also incorporates explicit height information, under comparable configurations. On the testing set, we evaluated our method using two mainstream backbone networks. With VoVNet-99 as the backbone, our approach achieved 50.5% mAP and 59.8% NDS. Notably, our model demonstrated exceptional performance on the mAAE metric, likely due to its comprehensive consideration of spatial correlations of an object. Tab. IV reports the performance of our $base$ configuration on the validation set of the Lyft dataset. Our method consistently surpasses the baseline model and significantly outperforms several LiDAR-based 3D object detection approaches. Fig. 6 presents several robust detection results on Lyft validation set. Our model demonstrates exceptional performance in handling challenging scenarios such as large objects in cross-view and backlighting conditions on the Lyft dataset as well.
TABLE II: 3D object detection results on nuScenes $val$ set. *: Trained with Lidar data. †: Pre-trained from FCOS3D. ‡: Pre-trained from DD3D.
@@
TABLE III: 3D detection results on nuScenes $test$ set. *: Trained with Lidar data. †: Pre-trained from FCOS3D. ‡: Pre-trained from DD3D.
@@
BEVDet4D[11] $1600\times 640$ Swin-T 0.451 0.569 0.511 0.241 0.386 0.301 0.121
GraphDETR3D†[20] $1600\times 640$ ResNet101 0.418 0.472 0.668 0.250 0.440 0.876 0.139
PETRv2†[18] $1600\times 640$ ResNet101 0.456 0.553 0.601 0.249 0.391 0.382 0.123
HV-BEVbase†(Ours) $1600\times 900$ ResNet101 0.464 0.556 0.604 0.261 0.380 0.393 0.132
DETR3D‡[14] $1600\times 900$ VoVNet-99 0.412 0.479 0.641 0.255 0.394 0.845 0.133
BEVDepth*‡[12] $1408\times 512$ VoVNet-99 0.503 0.600 0.445 0.245 0.378 0.320 0.126
GraphDETR3D‡[20] $1600\times 640$ VoVNet-99 0.425 0.495 0.621 0.251 0.386 0.790 0.128
PETRv2‡[18] $1600\times 640$ VoVNet-99 0.490 0.582 0.561 0.243 0.361 0.343 0.120
VEDet‡[53] $1600\times 640$ VoVNet-99 0.505 0.585 0.545 0.244 0.346 0.421 0.123
HeightFormer‡[37] $1600\times 900$ VoVNet-99 0.481 0.573 - - - - -
HV-BEVbase‡(Ours) $1600\times 900$ VoVNet-99 0.505 0.598 0.544 0.249 0.353 0.318 0.117
TABLE IV: 3D detection results on Lyft $val$ set.
[…]
Second[58] Lidar FPN 0.101
PointPillars[59] Lidar FPN 0.127
SSN[60] Lidar SECFPN 0.121
BEVFormer camera ResNet101-FPN 0.136
HV-BEV(Ours) camera ResNet101-FPN 0.153
(a) Case 1
(b) Case 2
Figure 6: Robust visualization results on Lyft validation set. Different colors of 3D bounding boxes denote different object classes.
IV-D Occupancy Perception Results
In addition to 3D object detection, we further validate the transferability of our model on the 3D occupancy prediction task. Specifically, the 256-dimensional BEV features are evenly partitioned into 16 height bins, with each bin assigned a 16-dimensional feature representation. This process transforms the BEV features into 3D volume features with an explicit height dimension. A 3D encoder-decoder network[56] is then appended as a segmentation head to produce the final 3D semantic occupancy predictions. Moreover, we integrate our method into a BEVFormer-style 3D occupancy predictor, SurroundOcc[57]. Since SurroundOcc assigns a 3D reference point to each volume query, we embed the proposed DHCA module solely into the 2D-3D cross-view attention module of SurroundOcc to enhance spatial feature aggregation. The configuration used follows the base setup described in Sec. IV-B, where the backbone is a ResNet101-DCN pretrained on FCOS3D. Tab. V reports the 3D semantic occupancy results on the nuScenes validation set. For evaluation, we adopt the IoU of occupied voxels (ignoring semantic classes) as the metric for scene completion (SC), and the mean IoU across all semantic classes as the metric for semantic scene completion (SSC). Under the simple replacement of the task head, our model outperforms BEVFormer by 1.28% in SC IoU and 3.22% in SSC mIoU. Additionally, incorporating the DHCA module into SurroundOcc also yields a slight improvement in overall prediction accuracy, particularly for larger objects, indicating the effectiveness of our method in enhancing feature aggregation for large-scale structures. Several qualitative comparisons are shown in Fig. 7. It can be observed that, after integrating our proposed method, the occupancy prediction model yields more complete predictions for relatively larger obstacles and exhibits improved accuracy in semantic classification.
TABLE V: 3D semantic occupancy results on nuScenes $val$ set. All methods are trained with dense occupancy labels in [57].
Model SC IoU SSC mIoU barrier   bicycle   bus   car   const. veh.   motorcycle   pedestrian   traffic cone   trailer   truck   drive. surf.   other flat   sidewalk   terrain   manmade   vegetation
BEVFormer[15] 30.50 16.75 14.22 6.58 23.46 28.28 8.66 10.77 6.64 4.05 11.20 17.78 37.28 18.00 22.88 22.17 13.80 22.21
HV-BEV 31.78 19.97 21.03 10.54 28.42 30.60 10.67 14.86 9.05 8.38 13.98 23.61 40.09 22.83 24.12 24.30 14.69 22.27
SurroundOcc[57] 31.49 20.30 20.59 11.68 28.06 30.86 10.70 15.14 14.09 12.06 14.38 22.06 37.29 23.70 24.49 22.77 14.89 21.86
SurroundOcc-DHCA 32.93 20.97 22.14 11.02 29.57 31.82 10.68 15.26 13.30 11.30 13.89 23.81 40.40 23.19 26.37 24.69 15.30 22.76
(a) Case 1
(b) Case 2
Figure 7: A qualitative comparison of 3D semantic occupancy prediction results between SurroundOcc and SurroundOcc-DHCA on two samples from the nuScenes validation set. The bottom-left corner shows the predictions from SurroundOcc, while the bottom-right corner presents the results from SurroundOcc-DHCA. Colored elliptical boxes highlight regions where the prediction performance of the two methods differs significantly.
IV-E Ablation Study